OpenStudy (ohmybookness):

I solved everything else I just need to solve #6-10. it says to solve it algebraically & to use a graph to demonstrate that my answers are correct. I believe I already have a graph I just need help solving them algebraically.

OpenStudy (phi):

I would redraw the answer to #2 to show the circles touching only at the 1 point.

OpenStudy (ohmybookness):

Done. Do you think you can help me with #6-10?

OpenStudy (ohmybookness):

I think i have #6 completed.

OpenStudy (ohmybookness):

I think i have # 7 also. So just numbers 8-10.

OpenStudy (phi):

For #6, reorder the 2nd equation to make the x term first \[ 32x^2 +9y^2 =324\\ -x^2 +3y^2 = 3\] I see if the bottom y^2 term had a -9 in front of it (instead of 3) we could add the two equations and eliminate y^2. So I would multiply the bottom equation by -3 all terms , both sides: \[ 32x^2 +9y^2 =324\\ 3x^2 -9y^2 = -9\]

OpenStudy (ohmybookness):

Okay yay! I have #6 correct

OpenStudy (phi):

if we add the two equations we get \[ 35x^2 +0 = 315 \\ 35x^2= 315\] divide both sides by 35 \[ x^2= \frac{315}{35}= 9 \] take the square root of both sides \[ x = \pm \sqrt{9}= \pm 3 \] so x=-3 or x=3 we now need to find the y values

OpenStudy (phi):

can you find the y values using algebra? you would pick either equation (I would choose the 2nd one because it has smaller numbers) for x=-3 you will get two y values and ditto for x=3

OpenStudy (ohmybookness):

would it be 2?

OpenStudy (ohmybookness):

y=-2 and y=2

OpenStudy (phi):

using the 2nd equation, written this way \[ 3y^2 = 3+x^2\] if x is 3 we get \[ 3y^2= 3+9=12\\y^2=4\\y= \pm2\] so (-3,2) and (-3,-2) are solutions for x=+3 we will get the same two y values, so (3,2) and (3,-2) are the other two solutions

OpenStudy (ohmybookness):

Thank you!!

OpenStudy (ohmybookness):

Im not sure if i have the work right for #7

OpenStudy (phi):

for #7, do you have any thoughts ? I would multiply the first equation by -3 what do you get ?

OpenStudy (ohmybookness):

I converted the 2nd equation to \[21x^2 =15y^2-60y+9\]

OpenStudy (phi):

how did you get -60y ?

OpenStudy (ohmybookness):

Well I aslo converted the first one \[7x^2 = 5y^2-20y+3\]

OpenStudy (phi):

I see the first equation is \[ 7y^2 -5x^2+20x=3\] (no y term). multiply that equation by -3

OpenStudy (phi):

you should get \[ -21y^2 +15x^2 -60x = -9\]

OpenStudy (ohmybookness):

Yes

OpenStudy (phi):

can you add the two equations?

OpenStudy (ohmybookness):

im working on it on a separate piece of paper

OpenStudy (phi):

what do you get ?

OpenStudy (ohmybookness):

sorry. I'm really stuck on this one

OpenStudy (phi):

to add two equations you write down the left side and then a plus sign and then the left side of the other equation adding \[ -21y^2 +15x^2 -60x = -9 \\ 21y^2+5x^2=209\] you could write \[ -21y^2 +15x^2 -60x + 21y^2+5x^2 =-9+209\]

OpenStudy (phi):

you can simplify the right side, right ?

OpenStudy (phi):

on the left side, combine "like terms" notice you have \(-21y^2 +21y^2\) 21 y^2 take away 21 y^2 leaves no y^2 (they cancel out) you also have 15 x^2 plus 5 x^2 that can be combined to get 20x^2

OpenStudy (ohmybookness):

so it's y2 +3y2 so far

OpenStudy (ohmybookness):

Im really bad at this :/

OpenStudy (phi):

here is what we have when we add the two equations \[ -21y^2 +15x^2 -60x + 21y^2+5x^2 =-9+209 \] on the right side, what is -9+209 ?

OpenStudy (ohmybookness):

the sum?

OpenStudy (phi):

yes, what is -9+209 ?

OpenStudy (ohmybookness):

uh

OpenStudy (ohmybookness):

Oh my god

OpenStudy (ohmybookness):

im so stupid

OpenStudy (ohmybookness):

200

OpenStudy (phi):

yes, 200. so we have, so far, \[ -21y^2 +15x^2 -60x + 21y^2+5x^2 = 200\]

OpenStudy (phi):

next, notice you have \(-21y^2 +21y^2 \) do you know what that adds up to ?

OpenStudy (phi):

you have 21 y squared and you take away 21 y squared. how many y squared are left ?

OpenStudy (ohmybookness):

0 or 1

OpenStudy (ohmybookness):

would it just by y

OpenStudy (phi):

0 21y^2 - 21 y^2 is 0 so we can ignore those two terms, and we have \[ 15x^2 -60x +5x^2 = 200\]

OpenStudy (phi):

next, notice we have two terms with an x^2 we have 15 x squared plus another 5 x squared how many x squared do we have ?

OpenStudy (ohmybookness):

20x4 or 20x2

OpenStudy (phi):

20x^2 (you don't change the x^2 to x^4) you now have \[ 20x^2-60x= 200\]

OpenStudy (phi):

you can add -200 to both sides \[ 20x^2-60x-200 = 200-200\] or \[ 20x^2 -60x-200=0\] luckily we can simplify that because we can divide each term by 20 (or if you know how to factor, we can factor out 20, like this: \[ 20(x^2-3x-10)=0\]

OpenStudy (phi):

now divide both sides by 20 \[ \frac{20}{20} (x^2-3x-10)= \frac{0}{20} \\ x^2 -3x-10= 0\]

OpenStudy (phi):

so "solve for x" you should factor that equation do you know how ?

OpenStudy (ohmybookness):

no ;/

OpenStudy (ohmybookness):

would it be positive and negative 2

OpenStudy (phi):

it's a bit complicated to explain how to factor you would first look at the last number 10, and write all pairs that multiply to get 10. they are: 1,10 2,5 next, we see the sign on the 10 is -10 (minus). that means the factors are different signs the sign on the middle term -3x is -. that means the biggest factor is minus put in the signs and add +1-10= -9 +2-5= -3 if any add up to the middle number -3, those are the factors. in this case +2, and -5 the factors are (x+2)(x-5)=0

OpenStudy (phi):

we now use the idea that either x+2=0 or x-5 =0 and that means x=-2 or x= 5

OpenStudy (ohmybookness):

so would this be the answer? (-2,3) (-2,-3) (5,2) (5,-2)

OpenStudy (phi):

yes, all your answers are correct. But you have to show your algebra

OpenStudy (ohmybookness):

Thank you! Sorry i took so long..